A normal population has a mean of 50 and standard distribution of 10. Selecting 16 sample compute probability of sample mean



a) greater than 60



b) less than 45



c) between 45 and 60

1
Expert's answer
2022-05-04T15:04:07-0400

We have a normal distribution, "\u03bc=50,\u03c3=10,n=16."

Let's convert it to the standard normal distribution.

"\bar{z}=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}."


"\text{a) } \bar{z}_1=\cfrac{60-50}{10/\sqrt{16}}=4,\\
P(\bar{X}>60)=\\
=P(\bar{Z}>4)=\\
=1-P(\bar{Z}<4)=\\
=1-1=0\text{ (from z-table).}"


"\text{b) } 
\bar{z}_2=\cfrac{45-50}{10/\sqrt{16}}=-2,\\
P(\bar{X}<45)=\\
=P(\bar{Z}<-2)=\\
=0.0228\text{ (from z-table).}"


"\text{c) } 
P(45<\bar{X}<60)=\\
=P(-2<\bar{Z}<4)=\\
=P(\bar{Z}<4)-P(\bar{Z}<-2)=\\
=1-0.0228=0.9772."